One ox eats the same amount as 15 sheep. Two people rent a field for $175. Person A puts 3 oxen in the field. Person B puts 30 sheep in the field. What should each person’s part of the rent be, based on the amount of grass the animals eat?
a. Person A should pay $100 and Person B should pay $75.
b. Person A should pay $105 and Person B should pay $70.
c. Person A should pay $90 and Person B should pay $85.
d. Person A should pay $92.50 and Person B should pay $82.50.
e. Person A should pay $87.50 and Person B should pay $87.50

Answer by mananth(16946)

1 ox = 15 sheep in terms eating grass
3 ox = 45 sheepin terms of eating grass
..
A puts 3 ox (same as 45 sheep ) in the field
B puts 30 sheep in the field.
..
ratio of A to B interms of rent = 45/30 =3/2
there are 2+3 = 5 parts in 175
A will have to pay 3/5 *175 =$ 105
So B will pay $ 70